How many ways 4 letters can be selected from 
the word RANDHEER
(A)2 Identical +2 distinct 

Grade:12th pass

2 Answers

Rajat
3 years ago
There are 2 R's and 2 E's
So number of ways we can choose 2 identical letters = 2 (either pair of R and E)
Say we have chosen R pair
Thus we are left with A,N,D,H,E,E
1) Case 1: E is NOT chosen among the distinct letters. 4C2= 6
2) case 2: E is chosen among the distinct letters . 4C1= 4
So required number of ways= 2(6+4)= 20
Arun
3 years ago
Dear student
 
if only selection is to be made then ways to select 2 identical letter = 2C1
now 2 different letters can be selected in 5C2 = 10
 
hence total ways = 2 * 10 = 20